40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Jun 13 19:00:26 2012, Sarah, DailyProgress, Laser, Transfer Functions 12x
    Reply  Fri Jun 22 22:34:28 2012, tara, DailyProgress, NoiseBudget, RIN coupling to Frequency noise 
    Reply  Wed Jun 27 18:51:56 2012, Sarah, DailyProgress, Laser, Transfer Functions Slide1.pngtf_oa_calc_measured.pngTF_allfreq.pngTF_fit_allfreq.pngnb_allfreq.png
       Reply  Fri Jul 6 18:59:49 2012, tara, DailyProgress, Laser, Transfer Functions IMG_1447.jpgRIN_Fnoise.pngRIN_Fnoise.png
          Reply  Tue Jul 10 02:55:37 2012, tara, DailyProgress, Laser, Transfer Functions 
             Reply  Mon Jul 16 19:08:34 2012, tara, DailyProgress, NoiseBudget, Transfer Functions (RIN to Frequency noise via photothermal) 
                Reply  Tue Jul 17 19:05:32 2012, tara, DailyProgress, NoiseBudget, Transfer Functions (RIN to Frequency noise via photothermal) 7x
                   Reply  Tue Jul 24 17:02:35 2012, Sarah, DailyProgress, NoiseBudget, Transfer Functions (RIN to Frequency noise via photothermal) 8x
                      Reply  Fri Aug 3 02:46:15 2012, tara, DailyProgress, NoiseBudget, Transfer Functions (RIN to Frequency noise via photothermal) RIN_coupling_2012_08_03.pngRIN_coupling.fig
                         Reply  Tue Aug 13 21:45:51 2013, tara, DailyProgress, NoiseBudget, Transfer Functions (RIN to Frequency noise via photothermal) farsi_2013_08_13.pngfarsi_2013_08_13.figRIN_TO_algaas.pngRIN_TO_algaas.fig
                            Reply  Fri Aug 16 04:35:58 2013, tara, DailyProgress, NoiseBudget, Transfer Functions (RIN to Frequency noise via photothermal) RIN_req_algaas.pngRIN_req_algaas.fig
                            Reply  Thu Aug 22 13:36:19 2013, tara, DailyProgress, NoiseBudget, Transfer Functions (RIN to Frequency noise via photothermal) fasi_2013_08_22.pngFarsi_compare.fig
          Reply  Wed Jul 11 11:13:38 2012, Sarah, DailyProgress, Laser, Transfer Functions 7x
          Reply  Sun Jul 15 23:20:32 2012, tara, DailyProgress, NoiseBudget, Transfer Functions (power fluctuation to Frequency noise via photothermal effect) Farsi_compare.pngFarsi_compare.fig
             Reply  Thu Jul 19 03:01:26 2012, tara, DailyProgress, NoiseBudget, Transfer Functions (power fluctuation to Frequency noise via photothermal effect) Farsi_compare.pngFarsi_compare.fig
                Reply  Thu Jul 19 23:56:45 2012, rana, DailyProgress, NoiseBudget, Transfer Functions (power fluctuation to Frequency noise via photothermal effect) 
                   Reply  Fri Jul 20 01:32:45 2012, tara, DailyProgress, NoiseBudget, Transfer Functions (power fluctuation to Frequency noise via photothermal effect) 
Message ID: 1005     Entry time: Wed Jun 27 18:51:56 2012     In reply to: 987     Reply to this: 1014
Author: Sarah 
Type: DailyProgress 
Category: Laser 
Subject: Transfer Functions 

Tara and I performed additional measurements of the transfer function for the coupling between power fluctuations and frequency.

Setup:

This time we altered the setup and instead of attenuating the signal input to the ACAV servo, we added an optical attenuator before the pd. We attenuated before the pd this time because of the possibility of pd saturation in the previous setup. The setup we used today was:

Slide1.png

Measurement and Data Analysis:

Consistent with the last measurements, the response function for the transfer function was the AOM Feedback and the reference function was the ACAV power fluctuations. We used the sine swept measurement group in the SR785 to get the data. We checked the TF by changing the power from 2.2mW to .22mW, and the data remained the same. 

We calibrated the raw data from dB to the units of Hz/RIN using the following relationship:

TF [Hz/RIN] = 10(TF[dB]/20) *710[Hz/V] * 3.6[V] *2

 

where the factor of 2 accounts for the double passed AOM, the factor of 710 is the calibration for the Marconi, and the 3.6 is the DC level.

The results for the TF are plotted below:

TF_allfreq.png

For the magnitude, the data taken at different frequency intervals overlapped perfectly. For the phase, the data did not overlap perfectly, so the data was concatenated with a slight jump at 100Hz.

Application to Noise Budget: 

 Using the previously measured RIN, I estimated the frequency noise with the transfer function just measured. First I fit the transfer function as shown below:

TF_fit_allfreq.png

Next I applied the following relationship to obtain the estimated frequency noise, and plotted it with the noise budget:

Frequency Noise: RIN * 3.6 [V] * 2 * 710 [Hz/V] * 10(TF[dB]/20)

 nb_allfreq.png

Compared to the RIN induced frequency noise with the previously measured TF (refer to elog entry attached to this one), the noise is lower at higher frequencies, drops off with a slightly steeper slope in the center, and slightly higher at lower frequencies (this could be due to error in fit though). Note that this RIN induced noise is only an estimate, and the calculation may be slightly incorrect due to the fact that the DC level was slightly different (by ~ a factor of 2) for the measurement of the intensity noise (previous elog) and the current transfer function measurement.

Comparing the calculated TF from elog 989 (using RIN from previous setup), the following plot is obtained:

tf_oa_calc_measured.png

Thus, the transfer function still does not match the calculated curve perfectly, but its a better match shapewise than the previous TF. Additionally, there still might be an error in the calculated curve.

 

Conclusion:

 The TF looks considerably better than before, perhaps because we optimized the phase for the EAOM two days ago, so we are driving the amplitude with more modulation.

 

== Note on the measurement ==

  •       RCAV is locked with 100 uW input power, Common/Fast gains on the TTFSS are 950/750. The reason to use low power is that we want to make sure that the effect due to power modulation will not be common in both paths. However, we need enough power so that RCAV loop gain is still high enough so that the frequency noise in the laser is suppressed and does not cover the RIN induced noise in ACAV loop.
  •       ACAV input power is 2.2 mW. To avoid having too much gain and unstable loop, we use an ND filter to reduce the power on RFPD down to the level it usually is at 1mW power (~200mV).
  •       To check if the TF is real or not, we tried to lower the power down to 0.2 mW and see no change in the measured TF. This is good, since the TF should be independent of the power input.
  •       The DC level from PDA100A @gain 20 is 3.6 V (the reference signal of our TF measurement). The modulation is 380mV pk-pk in sine wave form, so the TRANSPD_DC is ~ 3.6 +/- 0.19 V. We make sure that the PD is not saturated.
  •  ***** The measurement is made possible because the drift is very small. For the past 5 Hrs, it has been less than 1kHz in 5 minutes. **
ELOG V3.1.3-